### v1

# Minimality of free-boundary axial hyperplanes in high dimensional circular cones via calibration

Abstract.
Consider an $(n+1)$-dimensional circular cone. Using a calibration argument, we prove that if $n\geq 4$ and the aperture of the cone is sufficiently large, the intersection of the cone with an axial hyperplane is area-minimizing with respect to free-boundary variations inside the cone.
Key words and phrases:
relative perimeter, calibration, free boundary, minimal cones
2020 Mathematics Subject Classification:
49Q05, 49Q10, 35R35
The author thanks Gian Paolo Leonardi for constructive discussions on the topics of the paper. The author has been supported by GNAMPA (INdAM) Project 2023: “Esistenza e proprietà fini di forme ottime”; GNAMPA (INdAM) Project 2025: “Structures of sub-Riemannian hypersurfaces in Heisenberg groups”.
1. Introduction
Consider a volume $v>0$ of liquid inside a container, here identified with a Lipschitz open set $\Omega\subset\mathbb{R}^{3}$. According to Gauss’ formulation, in the absence of gravity, the configuration $E$ assumed by the liquid minimizes the free-energy functional
$P(F;\Omega)+\gamma\,P(F;\partial\Omega)$ (1.1)
among all sets $F\subset\Omega$ such that $P(F;\Omega)<+\infty$ and satisfying the volume constraint $|F|=v$. Here, $\gamma\in[-1,1]$ is the so-called relative adhesion coefficient between the liquid and $\partial\Omega$, and $P(F;U)$ stands for the relative perimeter of $F$ in a Borel set $U$ (see [26, 11] for more details).
From a mathematical point of view, given an open set $\Omega\subset\mathbb{R}^{n+1}$, a well-established internal regularity theory for minimizers of (1.1) is available in the literature. In fact, let $E$ be a minimizer of (1.1) and indicate by $\partial^{\ast}E$ its reduced boundary. Letting $M:=\overline{\partial^{\ast}E\cap\Omega}$, one can show that $M\cap\Omega$, that is, the internal portion of $M$, is formed by the union of a smooth hypersurface and a singular set $\Sigma$ whose Hausdorff dimension $\mathop{\rm dim}_{{\mathcal{H}}}\Sigma$ does not exceed the critical value n7n-7, being in particular nonempty only when $n\geq 7$; see [6] and the references therein.
Much less is known regarding the boundary regularity for minimizers of (1.1), that is, the regularity of the free-boundary interface $M\cap\partial\Omega$. One of the first general regularity results, which holds for smooth containers $\Omega\subset\mathbb{R}^{3}$, is due to Taylor [35], who proved that $M\cap\partial\Omega$ does not exhibit singularities. This result has been generalized much later by De Philippis and Maggi in [6] for minimizers of anisotropic capillarity functionals. In the isotropic case, which includes the energy functional (1.1), their result ensures that, if $\partial\Omega$ is $C^{1,1}$, then $M$ is the union of a $C^{1,1/2}$ hypersurface with boundary and a singular set $\Sigma$ with the property that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$. Further improvements on the Hausdorff dimension of $\Sigma\cap\partial\Omega$ have been provided in [4] when $\gamma$ is close to either 0, 11 or 1-1. See also [32] for stability and rigidity properties of singular capillary cones arising in the capillarity setting.
Focusing now on the special case $\gamma=0$, corresponding to the relative perimeter in $\Omega$, something stronger regarding the regularity of $M\cap\partial\Omega$ was proved by Grüter-Jost [15] and Grüter [14, 13] in the setting of varifolds, under the assumption that $\partial\Omega$ is $C^{2}$. In particular, in [15], the authors demonstrate that if a varifold $M$ is close to a free-boundary hyperplane in $\Omega$, then $M$ is $C^{1,\beta}$-graphical with respect to this hyperplane. Subsequently, in [13], the author exploits the latter result to show that, as in the interior case, the Hausdorff dimension of the singular set of the free-boundary satisfies $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-7$.
The common denominator of the boundary regularity results mentioned above, both for $\gamma\neq 0$ and $\gamma=0$, is that the container $\Omega$ must be smooth. The literature concerning boundary regularity becomes sparser when one looks for results involving nonsmooth containers $\Omega$. In this situation, the main available studies typically consider specific nonsmooth domains or classes of domains. This is the case, for example, in the series of works [16, 17, 18, 19] by Hildebrandt and Sauvigny, where the authors perform a thorough study of minimal surfaces in wedges (i.e., three-dimensional domains obtained as the intersection of two half-spaces); in particular, in [19, p. 73], they prove a regularity result for free-boundary minimal graphs on a plane orthogonal to the wedge.
The latter result has been recently generalized by Edelen and Li in [8]. In that work, they prove a boundary regularity theorem in the same spirit as Grüter-Jost [15], holding for a varifold $M$ under the assumption that $\Omega$ is locally polyhedral and that $M$ is close to a horizontal plane (for instance, if $\Omega=W^{2}\times\mathbb{R}^{n-1}$ is a wedge, horizontal planes are those containing $W^{2}\times\{0\}$). In the same work, Edelen and Li also prove that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-2$, with a sharper estimate holding when the dihedral angles of their cone models are nonobtuse; see [8, Theorem 1.2]. In the setting of their polyhedral domains, the singular set $\Sigma$ is defined as the complement of the points around which the interface $M$ is the graph of a $C^{1,\beta}$ function over a horizontal plane.
In this paper, we prove that in dimensions $n\geq 4$ and for circular cones with sufficiently large aperture, axial hyperplanes are area-minimizing with respect to free-boundary variations. This result fits naturally into some very recent contributions by the author in collaboration with Leonardi [25, 23]. In the latter, we establish a boundary monotonicity formula that holds for almost minimizers of the relative perimeter under minimal regularity assumptions on $\partial\Omega$, and we then employ this formula to prove a minimizing-cone property for a blow-up sequence at a boundary point of $\Omega$. In [25], the authors prove a Vertex-skipping theorem, which states that the boundary of any almost minimizer of the relative perimeter $E$ in a convex domain $\Omega\subset\mathbb{R}^{3}$ cannot contain isolated singularities (say, vertices) of $\partial\Omega$. We highlight that, after a dimension-reduction argument, this result combined with [8, Theorem 1.1] implies the bound $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$, improving the estimate given in [8, Theorem 1.2 (1)]. The study of the behavior of $E$ near a singularity of the container is also relevant from the perspective of free-boundary minimal surfaces, since the interface $M=\overline{\partial^{\ast}E\cap\Omega}$ is a free-boundary minimal hypersurface in $\Omega$. We refer to [3] and the references therein for background on free-boundary minimal surfaces.
In this framework, the main result proved here provides a counterexample to the validity of Vertex-skipping results in higher dimensions. The question of the validity of the Vertex-skipping for $n\geq 3$ was already tackled in [24]. In that work, using a suitable Lipschitz flow of deformations for a free-boundary surface in a Lipschitz container $\Omega$, which in particular allows $\partial\Omega$ (including its singular set) to move, we establish a stability result for axial hyperplanes in circular cones, stated as follows. Given an integer $n\geq 2$, and for $\lambda>0$, consider the circular cone
$\Omega_{\lambda}:=\left\{(x,t)\in\mathbb{R}^{n+1}\,:\,t>\lambda\sqrt{x_{1}^{2}+\dots+x_{n}^{2}}\,\right\}.$
Note that the opening angle $\alpha_{\lambda}\in(0,\pi)$ of $\Omega_{\lambda}$ is related with $\lambda$ by $\lambda=\cot(\alpha_{\lambda}/2)$. Then, if $n=2$, a hyperplane of $\mathbb{R}^{3}$ containing the axis of the cone (up to rotations, $x_{1}=0$) is unstable in $\Omega_{\lambda}$ for every $\lambda>0$. This is in accordance with the aforementioned Vertex-skipping theorem. In partial contrast, when $n\geq 3$, we show that there exists a threshold $\lambda^{\ast}(n)>0$ such that if $0<\lambda\leq\lambda^{\ast}(n)$, i.e., the circular cone has a sufficiently large opening angle $\alpha\geq\alpha_{\lambda^{\star}(n)}$, the axial hyperplane is stable in $\Omega_{\lambda}$, whereas for $\lambda>\lambda^{\ast}(n)$ it is unstable. In this paper, we intend to analyze whether, for $n\geq 3$ and $\lambda\leq\lambda^{\ast}(n)$, an axial hyperplane is merely stable or even area-minimizing in $\Omega_{\lambda}$. We will provide a basically positive answer for $n\geq 4$.
We say that a measurable set $E\subset\mathbb{R}^{n+1}$ has locally finite perimeter provided $P(E;U)<+\infty$ for every $U\subset\!\subset\mathbb{R}^{n+1}$, see [12, 26]. For $(x,t)\in\mathbb{R}^{n+1}$ and $r>0$, we denote by $B_{r}(x,t)$ the ball of radius $r$ centered at $(x,t)$ (with $B_{r}:=B_{r}(0,0)$). Our main result is the following:
Theorem 1.1.
For $n\geq 4$, consider the threshold parameter
$\bar{\lambda}(n):=\dfrac{1}{2}\dfrac{n-3}{\sqrt{n-2}},$ (1.2)
and assume that $0<\lambda\leq\bar{\lambda}(n)$. Let also
$E:=\{(x,t)\in\mathbb{R}^{n+1}\,:\,x_{1}>0\}.$
Then $E$ is a minimizer of the relative perimeter in $\Omega_{\lambda}$, that is, given a locally-finite perimeter set $F\subset\mathbb{R}^{n+1}$ satisfying $E\Delta F\subset\!\subset B_{R}$, for some $R>0$, one has
$P(E;\Omega_{\lambda}\cap B_{R})\leq P(F;\Omega_{\lambda}\cap B_{R}).$
We emphasize that the above theorem provides a definitive counterexample to the validity of a Vertex-skipping theorem such as [25, Theorem 1.1] in higher dimensions, but only when $n\geq 4$. The case $n=3$ is a limiting situation where, in view of [24], the portion of hyperplane $H_{\lambda}:=\partial E\cap\Omega_{\lambda}$ is stable with respect to free-boundary variations in $\Omega_{\lambda}$, $\lambda\geq\lambda^{\ast}$, but our minimality result, Theorem 1.1, does not apply. We also remark that, despite $\Omega_{\lambda}$ not being a polyhedral cone, Theorem 1.1 appears to suggest that the dimensional bound for the singular set of the free-boundary proved in [8, Theorem 1.2 (1)] cannot be improved beyond the threshold n4n-4. In fact, when $n=4$ and $\lambda\leq 1/2\sqrt{2}$, the result above implies that $\partial E\cap\partial\Omega_{\lambda}=\{0\}$, and so $0\in\Sigma\cap\partial\Omega_{\lambda}$.
The proof of Theorem 1.1 is based on a calibration argument. The calibration method was originally introduced by Bombieri, De Giorgi, and Giusti in their celebrated work [2] to prove the minimality of Simon’s cone in $\mathbb{R}^{8}$. Since then, several authors have contributed by proposing new versions of the original method [21, 10, 29, 5, 7] or by developing various generalizations to different contexts, such as the Steiner tree problem [27, 28, 33] and the Mumford-Shah functional [1, 30].
Here, the calibration method is adapted to the free-boundary setting in the sense that we require the calibrating vector field $Z$ to be tangent to $\partial\Omega_{\lambda}$. This turns out to be crucial for carrying out the calibration argument, as it allows us to exclude the contribution coming from the area of $(E\Delta F)\cap\partial\Omega_{\lambda}$. We point out that the vector field $Z$ will not be defined in the entire closed cone $\overline{\Omega}_{\lambda}$, but only outside a $2$-dimensional subset of $\overline{\Omega}_{\lambda}$. This will force us to perform an approximation argument to apply the Divergence Theorem on $(E\Delta F)\cap\Omega_{\lambda}$, see Section 5.
The existence of the vector field $Z$ will be shown in Section 4, and its construction will be carried out in two steps. Let $S_{\lambda}:=\partial\Omega_{\lambda}\setminus\{0\}$, and denote by $e_{1},\dots,e_{n},e_{t}$ the canonical basis of $\mathbb{R}^{n+1}$. We begin by calibrating the (n1)(n-1)-dimensional surface $S^{0}_{\lambda}:=H_{\lambda}\cap S_{\lambda}$ in $S_{\lambda}$; that is, we build a divergence-free vector field $Y$, defined on a suitable subset $S_{\lambda}^{\prime}$ of $S_{\lambda}$, such that $Y|_{S^{0}_{\lambda}}=e_{1}$ and $|Y|\leq 1$. In a slightly different formulation via differential forms, this was essentially done by Morgan in [31], using ideas from Lawlor’s PhD thesis [20]. Second, we extend the calibration obtained in the previous step to (a suitable portion of) $\Omega_{\lambda}$ by vertical projection onto $S_{\lambda}$: we will show that, due to the structure of the metric on $S_{\lambda}$, the vector field $Z$ obtained in this way is, in particular, divergence-free in its domain.
1.1. Organization of the paper.
In Section 2 we provide some preliminary notation and facts concerning differentiable manifolds (especially Riemannian ones) and sets of finite perimeter. In Section 3 we introduce further notation and perform some basic computations needed in the subsequent parts of the paper. Then, Section 4 is devoted to the construction of the calibration $Z$, whose main properties are collected in Theorem 4.1. Finally, in Section 5 we present the calibration argument; namely, we show how the existence of the vector field $Z$ from Theorem 4.1 can be used to demonstrate Theorem 1.1.
2. Preliminary notions and basic notation
2.1. Manifolds and $k$-forms.
Given a $n$-dimensional manifold $M$, we denote by $TM$, $T^{\ast}M$ the tangent and cotangent bundles of $M$ respectively. For any $k\geq 1$, we define the bundle $\bigwedge^{k}T^{\ast}M$ as the disjoint union of the vector spaces $\bigwedge^{k}(T_{p}M)^{\ast}$ (and we observe that $\bigwedge^{1}T^{\ast}M=T^{\ast}M$). When $k=0$, we identify $\bigwedge^{0}T^{\ast}M$ with $C^{\infty}(M)$. A function $\omega:M\to\bigwedge^{k}T^{\ast}M$ with the property that $\omega(p)\in\bigwedge^{k}(T_{p}M)^{\ast}$, for all $p\in M$, is called $k$-form. Given a local chart $\varphi=(x_{1},\dots,x_{n})$ for $M$, we denote by $\partial_{i}$ and $dx_{i}$, for $1\leq i\leq n$, the local frames induced by $\varphi$ on $TM$ and $T^{\ast}M$ respectively. We introduce the following multilinear algebra notation: given a basis $u_{1},\dots,u_{n}$ of a $n$-dimensional vector space $V$ and a multi-index $\alpha=(\alpha_{1},\dots,\alpha_{k})\in\mathbb{N}^{k}$ with $1\leq\alpha_{1}<\dots<\alpha_{k}\leq n$, we set
$u_{\alpha}=u_{\alpha_{1}}\wedge\dots\wedge u_{\alpha_{k}}.$
Moreover, fixing $1\leq i<j\leq k$, we define
$\begin{split}&\,\,\,u^{i}_{\alpha}=u_{\alpha_{1}}\wedge\dots\wedge u_{\alpha_{i-1}}\wedge u_{\alpha_{i+1}}\wedge\dots\wedge u_{\alpha_{k}}\\ u^{i,j}_{\alpha}=u_{\alpha_{1}}&\wedge\dots\wedge u_{\alpha_{i-1}}\wedge u_{\alpha_{i+1}}\wedge\dots\wedge u_{\alpha_{j-1}}\wedge u_{\alpha_{j+1}}\wedge\dots\wedge u_{\alpha_{k}}.\end{split}$ (2.1)
Any $k$-form can be then represented by
$\omega=\sum_{\underset{1\leq\alpha_{1}<\dots<\alpha_{k}\leq n}{\alpha=(\alpha_{1},\dots,\alpha_{k})\in\mathbb{N}^{k}}}\omega_{\alpha}\,dx_{\alpha},$
for suitable scalar functions $\omega_{\alpha}:M\to\mathbb{R}$. We indicate by $\mathfrak{X}(M)$, $\mathcal{A}^{k}(M)$ the space of the vector fields on $M$ and the space of the $k$-forms on $M$, respectively.
2.2. Riemannian manifolds.
Let $(M,g)$ be a Riemannian manifold. Given two vectors $v,w\in T_{p}M$, we set $\langle v,w\rangle_{g}:=g_{p}(v,w)$, $|v|_{g}:=\sqrt{g_{p}(v,v)}$, with the convention that, if the subscript $g$ is omitted, we are referring to the Euclidean scalar product. The metric $g$ can be expressed in terms of the local frame for $TM$, by introducing the $\mathbb{R}$-valued functions
$g_{i,j}(p):=\langle\partial_{i}(p),\partial_{j}(p)\rangle_{g}\,,\qquad\text{for $1\leq i,j\leq n$.}$
With a little abuse, we denote by $g$ the symmetric, positive-definite matrix $g=(g_{i,j})_{1\leq i,j\leq n}$, while we indicate by $\tilde{g}=(\tilde{g}_{i,j})_{1\leq i,j\leq n}$ the inverse of $g$. The metric $g$ on $M$ can be extended to $\bigwedge^{k}T^{\ast}M$, for all $k\geq 1$. When $k=1$, we set
$\langle dx_{i},dx_{j}\rangle_{g}:=\tilde{g}_{i,j}\,,\qquad\text{for $1\leq i,j\leq n$,}$
and, when $k\geq 2$, given two multi-indices $\alpha=(\alpha_{1},\dots,\alpha_{k}),\beta=(\beta_{1},\dots,\beta_{k})\in\mathbb{N}^{k}$, we define
$\begin{split}\langle dx_{\alpha},dx_{\beta}\rangle_{g}&:=\det(\langle dx_{\alpha_{i}},dx_{\beta_{j}}\rangle_{g})_{1\leq i,j\leq k}\\ &=\det(\tilde{g}_{\alpha_{i},\beta_{j}})_{1\leq i,j\leq k}.\end{split}$ (2.2)
These definitions can be then extended to $\bigwedge^{k}T^{\ast}M$ by multi-linearity. By employing the metric structure of a Riemannian manifold, one can establish an identification between the tangent and the cotangent bundles. In local coordinates, this identification acts as follows: given a vector field $X=\sum_{i=1}^{n}X_{i}\partial_{i}\in TM$ one can consider the $1$-form $X^{\flat}$ defined by
$X^{\flat}=\sum_{i,j=1}^{n}g_{i,j}X_{i}\,dx_{j}\,;$
conversely, given a $1$-form $\omega=\sum_{i=1}^{n}\omega_{i}\,dx_{i}\in T^{\ast}M$, one defines the vector field
$\omega^{\#}=\sum_{i,j=1}^{n}\tilde{g}_{i,j}\,\omega_{i}\,\partial_{j}\,.$
Note that $|X^{\flat}|_{g}=|X|_{g}$, $|\omega^{\#}|_{g}=|\omega|_{g}$, and also $X^{\flat\#}=X$, $\omega^{\#\flat}=\omega$.
Let now $M$ and $N$ be Riemannian manifolds and consider a diffeomorphism $F:M\to N$. Given a vector field $X\in\mathfrak{X}(M)$, the push-forward of $X$ with respect to $F$ is a vector field on $N$ defined by
$F_{\ast}X(q)=dF_{F^{-1}(q)}(X(F^{-1}(q)))\,,\qquad\text{for all $q\in N$.}$
We denote by $\operatorname{div}^{M}X$ the divergence of a vector field $X$ on a Riemannian manifold $M$, and we omit the superscript $M$ when referring to the Euclidean divergence. One can easily verify that, if $F$ is an isometry,
$\operatorname{div}^{N}F_{\ast}X\,(q)=\operatorname{div}^{M}X\,(F^{-1}(q)),\qquad\text{for every $q\in N$.}$ (2.3)
We conclude by stating the following formula for the divergence on $M$: if $X$ is expressed in local coordinates by $X=\sum_{i=1}^{n}X_{i}\partial_{i}$, we have
$\operatorname{div}^{M}X=\dfrac{1}{\sqrt{\det g}}\,\sum_{i=1}^{n}\,\partial_{i}\left(\sqrt{\det g}\,X_{i}\right),$ (2.4)
where $\det g$ is the determinant of the matrix $g=(g_{i,j})_{1\leq i,j\leq n}$ introduced above.
2.3. The Hodge-$\star$ operator.
Consider an oriented Riemannian manifold $(M,g)$ and denote by $\nu$ its volume form (in local coordinates, $\nu=\sqrt{\det g}\,dx_{1}\wedge\dots\wedge dx_{n}$). For any $k\in\mathbb{N}$, the Hodge-$\star$ operator is a linear isometry $\star:\bigwedge^{k}T^{\ast}M\to\bigwedge^{n-k}T^{\ast}M$ which is uniquely defined by the condition
$\omega\wedge(\star\psi)=\langle\omega,\psi\rangle_{g}\,\nu,\qquad\text{for all $k$-forms $\omega$ and $\psi$.}$ (2.5)
It can be shown that $\star\circ\star=\star^{2}$ coincides with the multiplication by $(-1)^{k(n-k)}$, and that $\star\nu=1$. In particular, when k=n1k=n-1, the operator $\star$ allows to associate to every (n1)(n-1)-form $\omega\in\mathcal{A}^{n-1}(M)$ a $1$-form $\star\omega\in\mathcal{A}^{1}(M)$. Given a vector field $X\in\mathfrak{X}(M)$, denote by $i_{X}\nu\in\mathcal{A}^{n-1}(M)$ the contraction between $X$ and the volume form $\nu$. We have the following technical Lemma (see [22, p. 52]).
Lemma 2.1.
Let $X\in\mathfrak{X}(M)$. The following identities hold:
  • (i)
    $\star X^{\flat}=i_{X}\nu$;
  • (ii)
    $\star d(i_{X}\nu)=\operatorname{div}^{M}X$.
2.4. Locally finite perimeter sets under set operations
Given a measurable set $E\subset\mathbb{R}^{n+1}$ and $s\in[0,1]$, we define
$E^{(s)}:=\left\{(x,t)\in\mathbb{R}^{n+1}\,:\,\lim_{\rho\to 0^{+}}\dfrac{|E\cap B_{\rho}(x,t)|}{|B_{\rho}(x,t)|}=s\right\}.$
If $E$ is a locally finite perimeter set, we indicate by $\partial^{\ast}E$ its reduced boundary and by $\nu_{E}$ the inner unit normal to its reduced boundary [26]. Assume now that $E$ and $F$ have locally finite perimeter. We denote by $Q(E,F)^{\pm}$ the set
$Q(E,F)^{\pm}:=\{(x,t)\in\partial^{\ast}E\cap\partial^{\ast}F\,:\,\nu_{E}(x,t)=\pm\nu_{F}(x,t)\}.$
It is well known that also $E\cap F$ and $E\setminus F$ have locally finite perimeter and, up to ${\mathcal{H}}^{n}$-negligible sets, one has (see [26, Theorem 16.3])
$\partial^{\ast}(E\cap F)=\left(F^{(1)}\cap\partial^{\ast}E\right)\cup\left(E^{(1)}\cap\partial^{\ast}F\right)\cup Q(E,F)^{+}$ (2.6)
$\partial^{\ast}(E\setminus F)=\left(F^{(0)}\cap\partial^{\ast}E\right)\cup\left(E^{(1)}\cap\partial^{\ast}F\right)\cup Q(E,F)^{-},$ (2.7)
and
[…]
3. Further notation and preliminary computations
Let $n\geq 2$ be an integer number. For $x^{\prime}\in\mathbb{R}^{n-1}$, we set $x=(x_{1},x^{\prime})$, and
$r:=|x^{\prime}|=\sqrt{x_{2}^{2}+\dots+x_{n}^{2}}\qquad\text{and}\qquad\rho:=|x|=\sqrt{x_{1}^{2}+r^{2}}\,.$
With this notation, the definition of $\Omega_{\lambda}$ given in the Introduction becomes
$\Omega_{\lambda}=\{(x,t)\in\mathbb{R}^{n+1}\,:\,t>\lambda\,\rho\}.$
We recall that $S_{\lambda}=\partial\Omega_{\lambda}\setminus\{0\}$ and we consider the following subsets of $\Omega_{\lambda}$ and $S_{\lambda}$, respectively:
$\Omega_{\lambda}^{\prime}:=\{(x,t)\in\Omega_{\lambda}\,:\,x^{\prime}\neq 0\}\qquad\text{and}\qquad S_{\lambda}^{\prime}:=\{(x,t)\in S_{\lambda}\,:\,x^{\prime}\neq 0\}.$
In other words, $\Omega_{\lambda}^{\prime}$ (resp. $S_{\lambda}^{\prime}$) coincides with $\Omega_{\lambda}$ (resp. $S_{\lambda}$) minus the $2$-plane $x^{\prime}=0$. We observe that the hypersurface $S_{\lambda}$ endowed with the Euclidean metric is a smooth Riemannian manifold. We consider the Riemannian manifold $M:=\mathbb{R}^{n}\setminus\{0\}$ equipped with the metric
$g_{i,j}(x):=\delta_{i,j}+\lambda^{2}\,\dfrac{x_{i}\,x_{j}}{\rho^{2}},\qquad\text{for $x\in M$ and $1\leq i,j\leq n$.}$ (3.1)
A direct computation yields the expression of the metric on the cotangent bundle
$\tilde{g}_{i,j}(x)=\delta_{i,j}-\dfrac{\lambda^{2}}{1+\lambda^{2}}\,\dfrac{x_{i}\,x_{j}}{\rho^{2}},\qquad\text{for $x\in M$ and $1\leq i,j\leq n$.}$
One can immediately check that $\mathcal{I}:M\to S_{\lambda}$, $\mathcal{I}(x):=(x,\lambda|x|)$, establishes an isometry between $S_{\lambda}$ and $M$. Moreover, $\mathcal{I}$ restricts to an isometry between $M^{0}:=\{x\in M\,:\,x_{1}=0\}$ and $S^{0}_{\lambda}:=S_{\lambda}\cap H_{\lambda}$. Consider the multi-index $\alpha=(2,\dots,n)$. Recalling notation (2.1), we introduce the (n2)(n-2)-form $\psi_{0}\in\mathcal{A}^{n-2}(M)$ defined by
$\psi_{0}(x):=\dfrac{\sqrt{1+\lambda^{2}}}{n-1}\,\sum_{i=2}^{n}\,(-1)^{i}\,x_{i}\,dx_{\alpha}^{i},\qquad\text{for $x\in M$.}$ (3.2)
Additionally, denoting by $M^{\prime}:=\{x\in M\,:\,x^{\prime}\neq 0\}$, we let
$\theta:=\arctan(x_{1}/r)\in(-\pi/2,\pi/2),\qquad\text{for $x\in M^{\prime}$.}$ (3.3)
In the following technical result we gather some useful identities.
Lemma 3.1.
The following identities hold:
$\hskip 73.97733pt\det g=1+\lambda^{2}$ (3.4)
$|d\rho|_{g}=\dfrac{1}{\sqrt{1+\lambda^{2}}}\qquad|d\theta|_{g}=\dfrac{1}{\rho}\qquad\langle d\rho,d\theta\rangle_{g}=0.$ (3.5)
$\hskip 56.9055pt|\psi_{0}|_{g}=\dfrac{r}{n-1}\sqrt{1+\lambda^{2}}.$ (3.6)
Proof.
Fix $x\in M$, and, for $1\leq i\leq n$, let
$\partial_{i}=e_{i}+A_{i}\,e_{t}\qquad\text{where}\qquad A_{i}:=\lambda\,\dfrac{x_{i}}{|x|}\,.$
Standard properties of the determinant guarantee that $\det g=|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}$. Define the multi-index $\beta=(1,\dots,n-1)$. We have
$\partial_{1}\wedge\dots\wedge\partial_{n}$ $=e_{\beta}\wedge e_{n}+A_{n}\,e_{\beta}\wedge e_{t}+\sum_{i=1}^{n-1}(-1)^{n-1-i}A_{i}\,e^{i}_{\beta}\wedge e_{t}\wedge e_{n}.$ (3.7)
Now the wedge products appearing in (3.7) are orthonormal each other, hence
$|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}=1+\sum_{i=1}^{n}A_{i}^{2}=1+\lambda^{2},$
and this proves (3.4).
Let us now show the validity of (3.5). We start observing that, by the definitions of $\rho$ and $\theta$,
$d\rho=\sum_{i=1}^{n}\dfrac{x_{i}}{\rho}\,dx_{i}\qquad\text{and}\qquad d\theta=\dfrac{1}{\rho^{2}}\left(r\,dx_{1}-\dfrac{x_{1}}{r}\sum_{i=2}^{n}x_{i}\,dx_{i}\right).$ (3.8)
Hence we have
$\begin{split}|d\rho|_{g}^{2}&=\sum_{i,j=1}^{n}\dfrac{x_{i}x_{j}}{\rho^{2}}\,\langle dx_{i}\,,\,dx_{j}\rangle_{g}\\ &=\sum_{i,j=1}^{n}\dfrac{x_{i}x_{j}}{\rho^{2}}\left(\delta_{i,j}-\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{x_{i}x_{j}}{\rho^{2}}\right)\\ &=\dfrac{1}{\rho^{2}}\sum_{i=1}^{n}x_{i}^{2}-\dfrac{\lambda^{2}}{\rho^{4}(1+\lambda^{2})}\left(\sum_{i=1}^{n}x_{i}^{2}\right)\left(\sum_{j=1}^{n}x_{j}^{2}\right)\\ &=\dfrac{1}{1+\lambda^{2}};\end{split}$
$\begin{split}|d\theta|_{g}^{2}&=\dfrac{1}{\rho^{4}}\left(r^{2}|dx_{1}|_{g}^{2}-2x_{1}\sum_{i=2}^{n}x_{i}\,\langle dx_{i}\,,\,dx_{1}\rangle_{g}+\dfrac{x_{1}^{2}}{r^{2}}\sum_{i,j=2}^{n}x_{i}x_{j}\,\langle dx_{i}\,,\,dx_{j}\rangle_{g}\right)\\ &=\dfrac{1}{\rho^{4}}\left(r^{2}\left(1-\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{x_{1}^{2}}{\rho^{2}}\right)+2x_{1}\sum_{i=2}^{n}\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{x_{1}x_{i}^{2}}{\rho^{2}}+x_{1}^{2}\left(1-\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{r^{2}}{\rho^{2}}\right)\right)\\ &=\dfrac{1}{\rho^{4}}\left(r^{2}+x_{1}^{2}\right)\\ &=\dfrac{1}{\rho^{2}};\end{split}$
$\begin{split}\langle d\rho\,,\,d\theta\rangle_{g}&=\dfrac{1}{\rho^{2}}\left\langle\sum_{i=1}^{n}\dfrac{x_{i}}{\rho}\,dx_{i}\,,\,r\,dx_{1}-\dfrac{x_{1}}{r}\sum_{i=2}^{n}x_{i}\,dx_{i}\right\rangle_{g}\\ &=\dfrac{1}{\rho^{2}}\left(\dfrac{r}{\rho}\,x_{1}\,|dx_{1}|_{g}^{2}+\left(\dfrac{r}{\rho}-\dfrac{x_{1}^{2}}{r\rho}\right)\sum_{i=2}^{n}x_{i}\,\langle dx_{1}\,,\,dx_{i}\rangle_{g}-\dfrac{x_{1}}{r\rho}\sum_{i,j=2}^{n}x_{i}x_{j}\,\langle dx_{i}\,,\,dx_{j}\rangle_{g}\right)\\ &=\dfrac{1}{\rho^{2}}\left(\dfrac{r}{\rho}\,x_{1}\left(1-\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{x_{1}^{2}}{\rho^{2}}\right)-\dfrac{\lambda^{2}}{1+\lambda^{2}}\left(\dfrac{r}{\rho}-\dfrac{x_{1}^{2}}{r\rho}\right)\sum_{i=2}^{n}\dfrac{x_{1}x_{i}}{\rho^{2}}-\dfrac{x_{1}}{\rho}\,r\left(1-\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{r^{2}}{\rho^{2}}\right)\right)\\ &=\dfrac{1}{\rho^{2}}\left(\dfrac{r}{\rho}x_{1}\left(1-\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{x_{1}^{2}}{\rho^{2}}-\right)-\dfrac{\lambda^{2}}{1+\lambda^{2}}\left(\dfrac{r}{\rho}-\dfrac{x_{1}^{2}}{r\rho}\right)x_{1}\,\dfrac{r^{2}}{\rho^{2}}-\dfrac{x_{1}}{\rho}\,r\left(1-\dfrac{\lambda^{2}}{1+\lambda^{2}}\dfrac{r^{2}}{\rho^{2}}\right)\right)\\[5.69046pt] &=0.\end{split}$
Ultimately, we prove (3.6). We have
$\begin{split}|\psi_{0}|_{g}^{2}&=\langle\psi_{0},\psi_{0}\rangle_{g}\\ &=\dfrac{1+\lambda^{2}}{(n-1)^{2}}\left\langle\sum_{i=2}^{n}(-1)^{i}\,x_{i}\,dx_{\alpha}^{i}\,,\sum_{j=2}^{n}(-1)^{j}\,x_{j}\,dx_{\alpha}^{j}\right\rangle_{g}\\ &=\dfrac{1+\lambda^{2}}{(n-1)^{2}}\sum_{i,j=2}^{n}(-1)^{i+j}\,x_{i}\,x_{j}\,\langle dx_{\alpha}^{i}\,,dx_{\alpha}^{j}\rangle_{g}.\end{split}$ (3.9)
Denote by $\tilde{g}_{1}$ the $(n-1)\times(n-1)$ submatrix of $\tilde{g}$ obtained by cancelling the first row and the first column of $\tilde{g}$. By the definition of the metric on $\bigwedge^{k}T^{\ast}M$ given in (2.2) (in this case k=n2k=n-2) and the properties of the adjugate matrix (see [34, p. 73]), we have
$\begin{split}\langle dx_{\alpha}^{i}\,,dx_{\alpha}^{j}\rangle_{g}=(-1)^{i+j}\,(\tilde{g}_{1}^{-1})_{i,j}\,\det(\tilde{g}_{1}).\end{split}$ (3.10)
A mere computation yields
$(\tilde{g}_{1}^{-1})_{i,j}=\delta_{i,j}+\dfrac{\lambda^{2}}{\rho^{2}+\lambda^{2}\,x_{1}^{2}}\,x_{i}\,x_{j},$ (3.11)
The same calculation performed above to compute $\det g$ guarantees also that
$\det(\tilde{g}_{1})^{-1}=\det(\tilde{g}_{1}^{-1})=1+\dfrac{\lambda^{2}\,r^{2}}{\rho^{2}+\lambda^{2}\,x_{1}^{2}}.$ (3.12)
Consequently, inserting (3.12) and (3.11) inside (3.10), by (3.9), we infer that
$\begin{split}|\psi_{0}|_{g}^{2}&=\dfrac{1+\lambda^{2}}{n-1}\sum_{i,j=2}^{n}\,x_{i}\,x_{j}\,\left(\delta_{i,j}+\dfrac{\lambda^{2}}{\rho^{2}+\lambda^{2}\,x_{1}^{2}}\,x_{i}\,x_{j}\right)\left(1+\dfrac{\lambda^{2}\,r^{2}}{\rho^{2}+\lambda^{2}\,x_{1}^{2}}\right)^{-1}\\ &=\dfrac{1+\lambda^{2}}{n-1}\,r^{2},\end{split}$
that is precisely (3.6). ∎
4. Construction of the calibration
The main result of this section is the following Theorem.
Theorem 4.1.
For $n\geq 4$, consider the threshold $\bar{\lambda}(n)$ defined in (1.2). Then, if $0<\lambda\leq\bar{\lambda}(n)$, there exists a vector field $Z:\Omega_{\lambda}^{\prime}\cup S_{\lambda}^{\prime}\to\mathbb{R}^{n+1}$ such that:
  • (i)
    $Z\left(x,t\right)=Z\left(x,\lambda|x|\right)$, for all $(x,t)\in\Omega_{\lambda}^{\prime}\cup S_{\lambda}^{\prime}$;
  • (ii)
    $|Z(x,t)|\leq 1$, for all $(x,t)\in\Omega_{\lambda}^{\prime}\cup S_{\lambda}^{\prime}$;
  • (iii)
    $Z(x,t)=e_{1}$, for all $(x,t)\in\overline{H}_{\lambda}\cap(\Omega_{\lambda}^{\prime}\cup S_{\lambda}^{\prime})$;
  • (iv)
    $Z(x,t)\in T_{(x,t)}S_{\lambda}$, for all $(x,t)\in S_{\lambda}^{\prime}$;
  • (v)
    $Z$ is smooth and $\operatorname{div}Z=0$ in $\Omega_{\lambda}^{\prime}$.
A key passage for the proof of Theorem 4.1, is the construction of a calibration for the (n1)(n-1)-dimensional surface $S^{0}_{\lambda}$ inside $S_{\lambda}$. This is the subject of the next result.
Theorem 4.2.
For $n\geq 4$, let $0<\lambda\leq\bar{\lambda}(n)$. Then there exists a vector field $Y\in\mathfrak{X}(S_{\lambda}^{\prime})$ satisfying the following conditions:
  • (i)
    $|Y(p)|\leq 1$, for all $p\in S_{\lambda}^{\prime}$;
  • (ii)
    $Y(p)=e_{1}$, for all $p\in S^{0}_{\lambda}$;
  • (iii)
    $Y$ is smooth and $\operatorname{div}^{S_{\lambda}}Y=0$ on $S_{\lambda}^{\prime}$.
The proof of Theorem 4.2 can be in part found in [31]. For the sake of clarity, we provide a complete proof here.
Proof.
Let $\alpha$ be the multi-index $(2,\dots,n)$, and consider the (n1)(n-1)-form $\omega_{0}=\sqrt{1+\lambda^{2}}\,dx_{\alpha}$. A direct computation shows that, in $M^{\prime}$,
$\omega_{0}=d\psi_{0}=\dfrac{n-1}{r}\,dr\wedge\psi_{0},$ (4.1)
where $\psi_{0}$ is as in (3.2). Additionally, set $u:=x_{1}/r$. Given a differentiable function $h$, we define
$\psi_{h}(x):=h(u)\,\psi_{0}(x)\qquad\text{and}\qquad\omega_{h}(x):=d\psi_{h}(x),\qquad\text{for $x\in M^{\prime}$.}$
By the properties of the exterior derivative, we have
$\omega_{h}=dh\wedge\psi_{0}+h\,\omega_{0}\,$, (4.2)
where
$dh=\dfrac{h^{\prime}(u)}{r}\,dx_{1}-\dfrac{x_{1}}{r^{2}}\,h^{\prime}(u)\,dr\,.$ (4.3)
Putting together (4.2), (4.3) and using (4.1), we infer that
$\omega_{h}(x)$ $=\left(\dfrac{h^{\prime}(u)}{r}\,dx_{1}-\dfrac{x_{1}}{r^{2}}\,h^{\prime}(u)\,dr\right)\wedge\psi_{0}(x)+h(u)\,\omega_{0}(x)$
$=\left[\dfrac{r}{n-1}\left(\dfrac{h^{\prime}(u)}{r}\,dx_{1}-\dfrac{x_{1}}{r^{2}}\,h^{\prime}(u)\,dr\right)+h(u)\,dr\right]\wedge\left(\dfrac{n-1}{r}\,\psi_{0}(x)\right)$ (4.4)
$=\left[\left(h(u)-\dfrac{x_{1}}{(n-1)\,r}h^{\prime}(u)\right)dr+\dfrac{h^{\prime}(u)}{n-1}\,dx_{1}\right]\wedge\left(\dfrac{n-1}{r}\,\psi_{0}(x)\right)\,.$
Step 1. We look for a continuously differentiable function $h:\mathbb{R}\to\mathbb{R}$ such that:
$h(0)=1,\qquad h^{\prime}(0)=0\qquad\text{and}\qquad|\omega_{h}(x)|_{g}\leq 1\quad\text{for all $x\in M^{\prime}$.}$ (4.5)
Applying the Cauchy-Schwartz inequality for wedge products (see [9, pag. 32], and note that $1$-forms are always simple), by (4), we obtain
$|\omega_{h}(x)|_{g}\leq\left|\left(h(u)-\dfrac{x_{1}}{(n-1)r}h^{\prime}(u)\right)dr+\dfrac{h^{\prime}(u)}{n-1}dx_{1}\right|_{g}\cdot\,\,\left|\dfrac{n-1}{r}\psi_{0}(x)\right|_{g}\,.$
Thanks to (3.6), to have $|\omega_{h}|_{g}\leq 1$, it suffices to impose
$\left|\left(h(u)-\dfrac{x_{1}}{(n-1)\,r}h^{\prime}(u)\right)dr+\dfrac{h^{\prime}(u)}{n-1}\,dx_{1}\right|_{g}^{2}\leq\dfrac{1}{1+\lambda^{2}}\,.$ (4.6)
We now observe that, taking $\theta\in(-\pi/2,\pi/2)$ as in (3.3), one has
$u=\tan\theta,\qquad x_{1}=\rho\sin\theta\qquad\text{and}\qquad r=\rho\cos\theta.$
The identities (3.5) permit us to rewrite (4.6) in the following way:
$h(u)^{2}+(1+\lambda^{2})\left[\dfrac{(1+u^{2})\,h^{\prime}(u)}{n-1}-u\,h(u)\right]^{2}\leq 1+u^{2}\,.$ (4.7)
Thus we are reduced to find a solution $h$ to the Cauchy problem for the differential inequality (4.7) with initial conditions $h(0)=1$, $h^{\prime}(0)=0$. In order to solve (4.7), we apply the ansatz
$h(u)=\dfrac{\cos(\beta(\arctan u))}{\cos(\arctan u)}=\dfrac{\cos(\beta(\theta))}{\cos\theta}\,$,
where $\beta:(-\pi/2,\pi/2)\to(-\pi/2,\pi/2)$ is a continuously differentiable function with $\beta(0)=0$. By this approach, $h(0)=1$. Moreover, we have
$h^{\prime}(u)=\dfrac{-\sin(\beta(\theta))\beta^{\prime}(\theta)\cos\theta+\cos(\beta(\theta))\sin\theta}{(1+u^{2})\cos^{2}(\theta)}\,$, (4.8)
and then, in particular, $h^{\prime}(0)=0$. We can now rewrite the differential inequality (4.7) in terms of $\beta$ and $\theta$. Thanks to (4.8), we have
$\dfrac{(1+u^{2})\,h^{\prime}(u)}{n-1}-u\,h(u)$ $=\dfrac{-\sin(\beta(\theta))\beta^{\prime}(\theta)\cos\theta+\cos(\beta(\theta))\sin\theta}{(n-1)\cos^{2}(\theta)}-\tan(\theta)\dfrac{\cos(\beta(\theta))}{\cos\theta}$
$=\dfrac{1}{n-1}\left(-\beta^{\prime}(\theta)\dfrac{\sin(\beta(\theta))}{\cos\theta}+\dfrac{\tan\theta}{\tan(\beta(\theta))}\dfrac{\sin(\beta(\theta))}{\cos\theta}\right)-\dfrac{\tan\theta}{\tan(\beta(\theta))}\dfrac{\sin(\beta(\theta))}{\cos\theta}$
$=\dfrac{\sin(\beta(\theta))}{\cos\theta}\left(-\dfrac{\beta^{\prime}(\theta)}{n-1}+\dfrac{2-n}{n-1}\dfrac{\tan\theta}{\tan(\beta(\theta))}\right)\,.$
Thus, by the definition of $h$ and the identity $\cos^{-2}(\theta)=1+u^{2}$, (4.7) becomes
$\left|\beta^{\prime}(\theta)+(n-2)\dfrac{\tan\theta}{\tan(\beta(\theta))}\right|\leq\dfrac{n-1}{\sqrt{1+\lambda^{2}}},\qquad\text{for $\theta\in(-\pi/2,\pi/2)$.}$ (4.9)
We want to find a solution $\beta$ to the above differential inequality with initial condition $\beta(0)=0$. We claim that such a solution can be found among the functions $\beta_{\gamma}:(-\pi/2,\pi/2)\to(-\pi/2,\pi/2)$ defined by
$\beta_{\gamma}(\theta):=\mathop{\mathrm{sgn}}(\theta)\arccos\left(\cos^{1+\gamma}(\theta)\right)\qquad\text{for $\gamma>0$,}$
i.e., that there exists $\gamma>0$ such that $\beta_{\gamma}$ satisfies (4.9) and $\beta_{\gamma}(0)=0$. Note that, for any $\gamma>0$, $\beta_{\gamma}(0)=0$. Let us rewrite (4.9) for $\beta_{\gamma}$. We observe that $\beta_{\gamma}$ is differentiable, and
$\beta_{\gamma}^{\prime}(\theta)=\mathop{\mathrm{sgn}}(\theta)\dfrac{(1+\gamma)\cos^{\gamma}\theta\,\sin\theta}{\sqrt{1-\cos^{2(1+\gamma)}(\theta)}}\,.$
Moreover, we have
$\dfrac{1}{\tan(\beta_{\gamma}(\theta))}=\dfrac{\cos(\beta_{\gamma}(\theta))}{\mathop{\mathrm{sgn}}(\beta_{\gamma}(\theta))\sqrt{1-\cos^{2}(\beta_{\gamma}(\theta))}}=\dfrac{\mathop{\mathrm{sgn}}(\theta)\cos^{1+\gamma}(\theta)}{\sqrt{1-\cos^{2(1+\gamma)}(\theta)}},$
and consequently,
$\left|\beta^{\prime}(\theta)+(n-2)\dfrac{\tan\theta}{\tan(\beta(\theta))}\right|=(n-1+\gamma)\left|\dfrac{\tan\theta}{\tan(\beta_{\gamma}(\theta))}\right|.$
Hence, for this family of functions $\beta_{\gamma}(\theta)$, (4.9) transforms into
$\left(1+\dfrac{\gamma}{n-1}\right)^{2}\tan^{2}\theta\leq\dfrac{\tan^{2}(\beta_{\gamma}(\theta))}{1+\lambda^{2}}\,,\qquad\text{for $\theta\in(-\pi/2,\pi/2)$.}$ (4.10)
Let us set $z:=\cos^{-2}(\theta)$. In this way, $\cos^{-2}(\beta_{\gamma}(\theta))=z^{1+\gamma}$, and condition (4.10) becomes
$\left(1+\dfrac{\gamma}{n-1}\right)^{2}\leq\dfrac{1}{1+\lambda^{2}}\,\dfrac{z^{1+\gamma}-1}{z-1}\,,\qquad\text{for $z\geq 1$.}$ (4.11)
Let $w(z):=(z^{1+\gamma}-1)/(z-1)$. For $z\geq 1$, $w$ is monotonically increasing, hence it suffices to test condition (4.11) only as $z\to 1^{+}$. Since $\lim_{z\to 1^{+}}w(z)=1+\gamma$, we get the condition
$\left(1+\dfrac{\gamma}{n-1}\right)^{2}-\dfrac{1+\gamma}{1+\lambda^{2}}\,\leq\,0\,.$ (4.12)
In particular, (4.12) implies (4.11). We now observe that the function on the LHS of (4.12) has a global minimizer in
$\bar{\gamma}:=(n-1)\left(\dfrac{n-1}{2(1+\lambda^{2})}-1\right).$
Let $\gamma=\bar{\gamma}$. If (4.12) is satisfied, by construction, the function $\beta=\beta_{\gamma}$ solves (4.9) and $\beta(0)=0$. A mere computation gives that (4.12) holds for $\gamma=\bar{\gamma}$ if and only if
$n\geq 4\qquad\text{and}\qquad 0<\lambda\leq\bar{\lambda}(n):=\dfrac{1}{2}\dfrac{n-3}{\sqrt{n-2}}\,$, (4.13)
and this concludes the proof of Step 1. In particular, when (4.13) is fulfilled, the function $h$ for which (4.5) holds true is given by
$h(u)$ $=\cos^{\gamma}(\arctan u),\qquad\text{where}\qquad\gamma=(n-1)\left(\dfrac{n-1}{2(1+\lambda^{2})}-1\right).$
Notice that $h$ is smooth, and so $\omega_{h}\in\mathcal{A}^{n-1}(M)$ is smooth.
Step 2. For $x\in M^{\prime}$, define
$X=(-1)^{n-1}(\star\,\omega_{h})^{\#}\in\mathfrak{X}(M^{\prime})\qquad\text{and}\qquad Y=\mathcal{I}_{\ast}X\in\mathfrak{X}(S_{\lambda}^{\prime}).$
We now show that $Y$ satisfies (i), (ii) and (iii). By construction, $|\omega_{h}|_{g}\leq 1$, and since $\star$ and $\mathcal{I}$ are isometries, we trivially infer that $|Y(p)|\leq 1$, for all $p\in S_{\lambda}^{\prime}$, that is (i). Moreover, by (4.5) and (4), since $u=0$ as $x_{1}=0$ , we have $\omega_{h}=\omega_{0}=\sqrt{1+\lambda^{2}}\,dx_{2}\wedge\dots\wedge dx_{n}$ on $M^{0}$. It is easy to check that, by (2.5) (note that $\nu=\sqrt{\det g}\,dx_{1}\wedge\dots\wedge dx_{n}=\sqrt{1+\lambda^{2}}\,dx_{1}\wedge\dots\wedge dx_{n}$), the following equality holds
$\star\left(\sqrt{1+\lambda^{2}}\,dx_{2}\wedge\dots\wedge dx_{n}\right)=(-1)^{n-1}e_{1}^{\flat},$
hence $X(x)=e_{1}$, for every $x\in M^{0}$. By the definition of $\mathcal{I}$, we have $Y(p)=e_{1}$, for all $p\in S^{0}_{\lambda}$, that is precisely (ii). Finally, let us prove (iii). The smoothness of $Y$ immediately follows from that of $\omega_{h}$. Now, by (i) of Lemma 2.1, $\star X^{\flat}=i_{X}\nu$, hence we have
$i_{X}\nu=(-1)^{n-1}\star^{2}\omega_{h}=\omega_{h}.$
By (ii) of Lemma 2.1 and the definition of $\omega_{h}$, we deduce that
$\operatorname{div}^{M}X=\star d\omega_{h}=\star dd\psi_{h}=0,$
and then (iii) immediately follows from (2.3). ∎
Let us now proceed with the proof of Theorem 4.1. The idea is to extend the vector field $Y$ of Theorem 4.2 to $\Omega_{\lambda}^{\prime}$ by vertically projecting onto $S_{\lambda}^{\prime}$. This operation will produce a divergence-free vector field in $\Omega_{\lambda}^{\prime}$.
Proof (of Theorem 4.1)..
Consider the vector field
$Z(x,t):=Y(x,\lambda|x|)\,.$
Since $Y$ is defined on $S_{\lambda}^{\prime}$, it is clear that $Z$ is defined in $\Omega_{\lambda}^{\prime}$. By construction, (i) of Theorem 4.1 holds, while the statements (ii), (iii), (iv) immediately follow from the properties of $Y$. It remains to demonstrate (v). For sure, $Z$ is smooth because $Y$ is. We need to show that $\operatorname{div}Z=0$ in $\Omega_{\lambda}^{\prime}$. Consider the (non-orthonormal) frame $\partial_{1},\dots,\partial_{n},\partial_{t}$ for $T\Omega_{\lambda}^{\prime}$ defined by
$\partial_{i}(x,t)=\partial_{i}(x)=e_{i}+\lambda\,\dfrac{x_{i}}{|x|}\,e_{t}\qquad\text{for $1\leq i\leq n$}\qquad\text{and}\qquad\partial_{t}=e_{t}.$
In addition, let $\hat{g}$ the matrix whose entries are the scalar products $\langle\partial_{i},\partial_{j}\rangle$, $\langle\partial_{i},\partial_{t}\rangle$, for $1\leq i,j\leq n$. By (3.7), the determinant of $\hat{g}$ satisfies
$\det\hat{g}$ $=|\partial_{1}\wedge\dots\wedge\partial_{n}\wedge e_{t}|^{2}$
$=|e_{1}\wedge\dots\wedge e_{n}\wedge e_{t}|^{2}$
$=1,$
and, in particular, is constant. Represent now $Y$ and $Z$ in coordinates with respect to $\partial_{1},\dots,\partial_{n}$ and $\partial_{1},\dots,\partial_{n},\partial_{t}$ respectively. Then, for suitable $Y_{i}\in C^{\infty}(S_{\lambda}^{\prime})$, $Z_{i},Z_{t}\in C^{\infty}(\Omega_{\lambda}^{\prime})$,
$Y(x,\lambda|x|)=\sum_{i=1}^{n}Y_{i}(x,\lambda|x|)\partial_{i}(x)\qquad\text{and}\qquad Z(x,t)=\sum_{i=1}^{n}Z_{i}(x,t)\partial_{i}(x)+Z_{t}(x,t)\partial_{t},$
for all $x\in\mathbb{R}^{n}$ and $t\geq\lambda|x|$. By construction, we have
$Z_{i}(x,t)=Y_{i}(x,\lambda|x|)\quad\text{for any $1\leq i\leq n$}\qquad\text{and}\qquad Z_{t}\equiv 0.$
Consequently, since by (3.4) also $\det g=1+\lambda^{2}$ is constant, (2.4) ensures that
$\operatorname{div}Z(x,t)=\sum_{i=1}^{n}\partial_{i}Z_{i}(x,t)=\sum_{i=1}^{n}\partial_{i}Y_{i}(x,\lambda|x|)=\operatorname{div}^{S_{\lambda}^{\prime}}Y(x,\lambda|x|)=0.$
This closes the proof. ∎
5. Minimality of $E$
In this section, we demonstrate Theorem 1.1. The idea is to apply the Divergence Theorem to the vector field $Z$ in the intersection between $E\Delta F$ and a sequence of sets invading $\Omega_{\lambda}^{\prime}\subset\Omega_{\lambda}$. Subsequently, since $Z$ is tangent to $S_{\lambda}^{\prime}$, by an approximation argument, we will be able to infer the minimality condition of Theorem 1.1.
Proof (of Theorem 1.1)..
Up to rescaling, we can assume $R=1$. For $\varepsilon>0$, let
$A^{\varepsilon}:=\{(x,t)\in\mathbb{R}^{n+1}\,:\,|x^{\prime}|\leq\varepsilon\}\qquad\text{and}\qquad\Omega_{\lambda}^{\varepsilon}:=\Omega_{\lambda}\setminus A^{\varepsilon}.$
We observe that, for every $\varepsilon>0$, $\Omega_{\lambda}^{\varepsilon}\subset\Omega_{\lambda}^{\prime}$. When $0<\lambda\leq\bar{\lambda}(n)$, we can consider the vector field $Z:\Omega_{\lambda}^{\prime}\to\mathbb{R}^{n+1}$ of Theorem 4.1. By the Divergence Theorem (see [26]) and since $Z$ is divergence-free, we have
$0=\int_{(F\setminus E)\cap\Omega_{\lambda}^{\varepsilon}}\operatorname{div}Z=\int_{\partial^{\ast}((F\setminus E)\cap\Omega_{\lambda}^{\varepsilon})}\left\langle Z\,,\,\nu_{(F\setminus E)\cap\Omega_{\lambda}^{\varepsilon}}\right\rangle\,d{\mathcal{H}}^{n-1}.$ (5.1)
Owing to (2.6), there holds
$\partial^{\ast}((F\setminus E)\cap\Omega_{\lambda}^{\varepsilon})=\left(\Omega_{\lambda}^{\varepsilon}\cap\partial^{\ast}(F\setminus E)\right)\cup\left((F\setminus E)^{(1)}\cap\partial\Omega_{\lambda}^{\varepsilon}\right)\cup Q\left(F\setminus E,\Omega_{\lambda}^{\varepsilon}\right)^{+},$
and this in combination with (5.1) and (2.8) ensures that
$\begin{split}0=\int_{(F\setminus E)\cap\Omega_{\lambda}^{\varepsilon}}\operatorname{div}Z&=\int_{\Omega_{\lambda}^{\varepsilon}\cap\partial^{\ast}(F\setminus E)}\left\langle Z\,,\,\nu_{F\setminus E}\right\rangle\,d{\mathcal{H}}^{n-1}+\\ &\qquad\qquad+\int_{\left((F\setminus E)^{(1)}\cap\partial\Omega_{\lambda}^{\varepsilon}\right)\,\cup\,Q\left(F\setminus E,\Omega_{\lambda}^{\varepsilon}\right)^{+}}\left\langle Z\,,\,\nu_{\Omega_{\lambda}^{\varepsilon}}\right\rangle\,d{\mathcal{H}}^{n-1}\\ &=(I)+(II).\end{split}$ (5.2)
By (2.7), because $Z=e_{1}=\nu_{E}$ on $\partial^{\ast}E\cap\Omega_{\lambda}=H_{\lambda}$, one has
$\begin{split}(I)&=\int_{\Omega_{\lambda}^{\varepsilon}\cap E^{(0)}\cap\partial^{\ast}F}\left\langle Z\,,\,\nu_{F}\right\rangle\,d{\mathcal{H}}^{n-1}-\int_{\Omega_{\lambda}^{\varepsilon}\cap F^{(1)}\cap\partial^{\ast}E}\left\langle Z\,,\,\nu_{E}\right\rangle\,d{\mathcal{H}}^{n-1}-\int_{\Omega_{\lambda}^{\varepsilon}\,\cap\,Q(E,F)^{-}}\left\langle Z\,,\,\nu_{E}\right\rangle\,d{\mathcal{H}}^{n-1}\\[5.69046pt] &\leq P(F;E^{(0)}\cap\Omega_{\lambda}^{\varepsilon})-P(E;F^{(1)}\cap\Omega_{\lambda}^{\varepsilon})\\ &=P(F;\bar{E}^{c}\cap\Omega_{\lambda}^{\varepsilon})-P(E;F^{(1)}\cap\Omega_{\lambda}^{\varepsilon}).\end{split}$ (5.3)
Let us now estimate $(II)$. Because $E\Delta F\subset\!\subset B_{1}$, clearly $\partial^{\ast}(F\setminus E)\subset B_{1}$ and $(F\setminus E)^{(1)}\subset B_{1}$. Thus, by applying (2.7) to $\Omega_{\lambda}^{\varepsilon}=\Omega_{\lambda}\setminus A^{\varepsilon}$ and (2.9) to its normal, we have
$\begin{split}|(II)|&\leq\int_{B_{1}\cap\partial\Omega_{\lambda}^{\varepsilon}}\left|\left\langle Z\,,\,\nu_{\Omega_{\lambda}^{\varepsilon}}\right\rangle\right|\,d{\mathcal{H}}^{n-1}\\ &\leq\int_{B_{1}\cap A^{\varepsilon}\cap\partial\Omega_{\lambda}}\left|\left\langle Z\,,\,\nu_{\Omega_{\lambda}}\right\rangle\right|\,d{\mathcal{H}}^{n-1}+\int_{B_{1}\cap\Omega_{\lambda}\cap\partial A^{\varepsilon}}\left|\left\langle Z\,,\,\nu_{A^{\varepsilon}}\right\rangle\right|\,d{\mathcal{H}}^{n-1}.\end{split}$ (5.4)
Now, since $Z$ is tangent to $S_{\lambda}^{\prime}$, the first integral on the right hand side of (5.4) vanishes. This implies that
$|(II)|\leq{\mathcal{H}}^{n-1}(B_{1}\cap\partial A^{\varepsilon})=C(n)\,\varepsilon^{n-2}.$
This in combination with (5.2) and (5.3) yields
$P(E;F^{(1)}\cap\Omega_{\lambda}^{\varepsilon})\leq P(F;\bar{E}^{c}\cap\Omega_{\lambda}^{\varepsilon})+C(n)\,\varepsilon^{n-2}.$ (5.5)
By applying the Divergence Theorem in $(E\setminus F)\cap\Omega_{\lambda}^{\varepsilon}$ and arguing very similarly as above, we also obtain
$\begin{split}P(E;F^{(0)}\cap\Omega_{\lambda}^{\varepsilon})&\leq P(F;E^{(1)}\cap\Omega_{\lambda}^{\varepsilon})+C(n)\,\varepsilon^{n-2}\\ &=P(F;E\cap\Omega_{\lambda}^{\varepsilon})+C(n)\,\varepsilon^{n-2}\end{split}$ (5.6)
Summing up (5.5) and (5.6), noticing that
$P(E;F^{(1)}\setminus B_{1})=P(E;F^{(0)}\setminus B_{1})=0\qquad\text{and}\qquad P(F;E\setminus B_{1})=P(F;\bar{E}^{c}\setminus B_{1})=0,$
we deduce that
$P(E;(F^{(0)}\cup F^{(1)})\cap\Omega_{\lambda}^{\varepsilon}\cap B_{1})\leq P(F;(\bar{E}^{c}\cup E)\cap\Omega_{\lambda}^{\varepsilon}\cap B_{1})+C(n)\,\varepsilon^{n-2}.$
By applying Federer’s Theorem [26, Theorem 16.2], we then get
$P(E;\Omega_{\lambda}^{\varepsilon}\cap B_{1})\leq P(F;\Omega_{\lambda}^{\varepsilon}\cap B_{1})+C(n)\,\varepsilon^{n-2}.$
Finally, passing to the limit as $\varepsilon\to 0^{+}$ in the relation above, we obtain that
$P(E;\Omega_{\lambda}\cap B_{1})\leq P(F;\Omega_{\lambda}\cap B_{1}),$
and we conclude. ∎

### v2

@@
 Abstract.
-Consider an $(n+1)$-dimensional circular cone.
-Using a calibration argument, we prove that if $n\geq 4$ and the aperture of the cone is sufficiently large, the intersection of the cone with an axial hyperplane is area-minimizing with respect to free-boundary variations inside the cone.
+Consider an $(n+1)$-dimensional circular cone with opening angle $\alpha\in(0,\pi)$.
+Using a free-boundary adaptation of the classical calibration method, we prove that, for $n\geq 4$, there exists a threshold $\bar{\alpha}(n)\in(0,\pi)$ such that if $\alpha\geq\bar{\alpha}(n)$, that is, the cone is wide enough, the intersection of the cone with an axial hyperplane is area-minimizing with respect to free-boundary variations inside the cone.
+This provides a counterexample to a recent Vertex-skipping Theorem proved by the author in collaboration with G.P.
+Leonardi, at least for $n\geq 4$.
 Key words and phrases:
-relative perimeter, calibration, free boundary, minimal cones
+relative perimeter, calibration, free-boundary, minimal cones
 2020 Mathematics Subject Classification:
@@
 The author thanks Gian Paolo Leonardi for constructive discussions on the topics of the paper.
-The author has been supported by GNAMPA (INdAM) Project 2023: “Esistenza e proprietà fini di forme ottime”; GNAMPA (INdAM) Project 2025: “Structures of sub-Riemannian hypersurfaces in Heisenberg groups”.
+The author has been supported by GNAMPA (INdAM) Project 2025: “Structures of sub-Riemannian hypersurfaces in Heisenberg groups”; GNAMPA (INdAM) Project 2026: “Variational, Geometric, and Analytic Perspectives on Regularity”.
 1. Introduction
@@
 In fact, let $E$ be a minimizer of (1.1) and indicate by $\partial^{\ast}E$ its reduced boundary.
-Letting $M:=\overline{\partial^{\ast}E\cap\Omega}$, one can show that $M\cap\Omega$, that is, the internal portion of $M$, is formed by the union of a smooth hypersurface and a singular set $\Sigma$ whose Hausdorff dimension $\mathop{\rm dim}_{{\mathcal{H}}}\Sigma$ does not exceed the critical value n7n-7, being in particular nonempty only when $n\geq 7$; see [6] and the references therein.
+Letting $M:=\overline{\partial^{\ast}E\cap\Omega}$, it is known that $M\cap\Omega$, that is, the internal portion of $M$, is formed by the union of a smooth hypersurface and a singular set $\Sigma$ whose Hausdorff dimension $\mathop{\rm dim}_{{\mathcal{H}}}\Sigma$ does not exceed the critical value n7n-7, being in particular nonempty only when $n\geq 7$; see [7] and the references therein.
 Much less is known regarding the boundary regularity for minimizers of (1.1), that is, the regularity of the free-boundary interface $M\cap\partial\Omega$.
@@
 Further improvements on the Hausdorff dimension of $\Sigma\cap\partial\Omega$ have been provided in [5] when $\gamma$ is close to either 0, 11 or 1-1.
-See also [32] for stability and rigidity properties of singular capillary cones arising in the capillarity setting.
+See also [9] for a regularity result on $\Lambda$-minimizers of (1.1) under the assumption that $\Omega$ is a half-space and the wet region of competitors is constrained within an open subset of $\partial\Omega$; and [38] for stability and rigidity properties of singular capillary cones.
 Focusing now on the special case $\gamma=0$, corresponding to the relative perimeter in $\Omega$, something stronger regarding the regularity of $M\cap\partial\Omega$ was proved by Grüter-Jost [17] and Grüter [16, 15] in the setting of varifolds, under the assumption that $\partial\Omega$ is $C^{2}$.
@@
 In this situation, the main available studies typically consider specific nonsmooth domains or classes of domains.
-This is the case, for example, in the series of works [16, 17, 18, 19] by Hildebrandt and Sauvigny, where the authors perform a thorough study of minimal surfaces in wedges (i.e., three-dimensional domains obtained as the intersection of two half-spaces); in particular, in [19, p.
+This is the case, for example, in the series of works [18, 19, 20, 21] by Hildebrandt and Sauvigny, where the authors perform a thorough study of minimal surfaces in wedges (i.e., three-dimensional domains obtained as the intersection of two half-spaces); in particular, in [21, p.
 73], they prove a regularity result for free-boundary minimal graphs on a plane orthogonal to the wedge.
 The latter result has been recently generalized by Edelen and Li in [10].
-In that work, they prove a boundary regularity theorem in the same spirit as Grüter-Jost [15], holding for a varifold $M$ under the assumption that $\Omega$ is locally polyhedral and that $M$ is close to a horizontal plane (for instance, if $\Omega=W^{2}\times\mathbb{R}^{n-1}$ is a wedge, horizontal planes are those containing $W^{2}\times\{0\}$).
-In the same work, Edelen and Li also prove that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-2$, with a sharper estimate holding when the dihedral angles of their cone models are nonobtuse; see [8, Theorem 1.2].
-In the setting of their polyhedral domains, the singular set $\Sigma$ is defined as the complement of the points around which the interface $M$ is the graph of a $C^{1,\beta}$ function over a horizontal plane.
-In this paper, we prove that in dimensions $n\geq 4$ and for circular cones with sufficiently large aperture, axial hyperplanes are area-minimizing with respect to free-boundary variations.
-This result fits naturally into some very recent contributions by the author in collaboration with Leonardi [25, 23].
-In the latter, we establish a boundary monotonicity formula that holds for almost minimizers of the relative perimeter under minimal regularity assumptions on $\partial\Omega$, and we then employ this formula to prove a minimizing-cone property for a blow-up sequence at a boundary point of $\Omega$.
+In that work, they prove a boundary regularity theorem in the same spirit as Grüter-Jost [17], holding for a varifold $M$ under the assumption that $\Omega$ is locally polyhedral and that $M$ is close to a horizontal plane (for instance, if $\Omega=W^{2}\times\mathbb{R}$ is a wedge, horizontal planes are those containing $W^{2}\times\{0\}$).
+In the same work, Edelen and Li also prove that $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-2$, with a sharper estimate holding when the dihedral angles of their cone models are nonobtuse; see [10, Theorem 1.2].
+In the setting of their polyhedral domains, the singular set $\Sigma$ is defined as the complementary set of the points around which the interface $M$ is the graph of a $C^{1,\beta}$ function over a horizontal plane.
+In this paper, we prove that for $n\geq 4$ and for circular cones with sufficiently large aperture, axial hyperplanes are area-minimizing with respect to free-boundary variations.
+This result fits naturally into some very recent contributions by the author in collaboration with Leonardi [28, 29] concerning boundary regularity in nonsmooth domains.
+In the latter, we establish a boundary monotonicity formula that holds for almost minimizers of the relative perimeter under mild regularity assumptions on $\partial\Omega$, which is then employed to prove a minimizing-cone property for a blow-up sequence at a boundary point of $\Omega$.
 In [28], the authors prove a Vertex-skipping Theorem, which states that the boundary of any almost minimizer of the relative perimeter $E$ in a convex domain $\Omega\subset\mathbb{R}^{3}$ cannot contain isolated singularities (say, vertices) of $\partial\Omega$.
-We highlight that, after a dimension-reduction argument, this result combined with [8, Theorem 1.1] implies the bound $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$, improving the estimate given in [8, Theorem 1.2 (1)].
-The study of the behavior of $E$ near a singularity of the container is also relevant from the perspective of free-boundary minimal surfaces, since the interface $M=\overline{\partial^{\ast}E\cap\Omega}$ is a free-boundary minimal hypersurface in $\Omega$.
-We refer to [3] and the references therein for background on free-boundary minimal surfaces.
+We highlight that, after a dimension-reduction argument, this result combined with [10, Theorem 1.1] implies the bound $\mathop{\rm dim}_{{\mathcal{H}}}(\Sigma\cap\partial\Omega)\leq n-3$, improving the estimate given in [10, Theorem 1.2 (1)].
 In this framework, the main result proved here provides a counterexample to the validity of Vertex-skipping results in higher dimensions.
@@
 Note that the opening angle $\alpha_{\lambda}\in(0,\pi)$ of $\Omega_{\lambda}$ is related with $\lambda$ by $\lambda=\cot(\alpha_{\lambda}/2)$.
-Then, if $n=2$, a hyperplane of $\mathbb{R}^{3}$ containing the axis of the cone (up to rotations, $x_{1}=0$) is unstable in $\Omega_{\lambda}$ for every $\lambda>0$.
+Then, if $n=2$, a hyperplane of $\mathbb{R}^{3}$ containing the axis of the cone (up to rotations, $x_{1}=0$) is unstable in $\Omega_{\lambda}$, for every $\lambda>0$.
 This is in accordance with the aforementioned Vertex-skipping Theorem.
-In partial contrast, when $n\geq 3$, we show that there exists a threshold $\lambda^{\ast}(n)>0$ such that if $0<\lambda\leq\lambda^{\ast}(n)$, i.e., the circular cone has a sufficiently large opening angle $\alpha\geq\alpha_{\lambda^{\star}(n)}$, the axial hyperplane is stable in $\Omega_{\lambda}$, whereas for $\lambda>\lambda^{\ast}(n)$ it is unstable.
-In this paper, we intend to analyze whether, for $n\geq 3$ and $\lambda\leq\lambda^{\ast}(n)$, an axial hyperplane is merely stable or even area-minimizing in $\Omega_{\lambda}$.
+In partial contrast, when $n\geq 3$, we show that there exists a threshold $\lambda^{\ast}(n)>0$ such that if $0<\lambda\leq\lambda^{\ast}(n)$, i.e., the circular cone has a sufficiently large opening angle $\alpha\geq\alpha^{\ast}(n):=\alpha_{\lambda^{\star}(n)}$, the axial hyperplane is stable in $\Omega_{\lambda}$, whereas for $\lambda>\lambda^{\ast}(n)$, or equivalently if $\alpha<\alpha^{\ast}(n)$, it is unstable.
+In this paper, we intend to analyze whether, for $n\geq 3$ and $0<\lambda\leq\lambda^{\ast}(n)$, an axial hyperplane is merely stable or even area-minimizing in $\Omega_{\lambda}$.
 We will provide a basically positive answer for $n\geq 4$.
 We say that a measurable set $E\subset\mathbb{R}^{n+1}$ has locally finite perimeter provided $P(E;U)<+\infty$ for every $U\subset\!\subset\mathbb{R}^{n+1}$, see [14, 31].
-For $(x,t)\in\mathbb{R}^{n+1}$ and $r>0$, we denote by $B_{r}(x,t)$ the ball of radius $r$ centered at $(x,t)$ (with $B_{r}:=B_{r}(0,0)$).
+For $(x,t)\in\mathbb{R}^{n+1}$ and $r>0$, we denote by $B_{r}(x,t)$ the ball of radius $r$ centered at $(x,t)$ (setting $B_{r}:=B_{r}(0,0)$).
 Our main result is the following:
@@
 Then $E$ is a minimizer of the relative perimeter in $\Omega_{\lambda}$, that is, given a locally-finite perimeter set $F\subset\mathbb{R}^{n+1}$ satisfying $E\Delta F\subset\!\subset B_{R}$, for some $R>0$, one has
-$P(E;\Omega_{\lambda}\cap B_{R})\leq P(F;\Omega_{\lambda}\cap B_{R}).$
-We emphasize that the above theorem provides a definitive counterexample to the validity of a Vertex-skipping theorem such as [25, Theorem 1.1] in higher dimensions, but only when $n\geq 4$.
-The case $n=3$ is a limiting situation where, in view of [24], the portion of hyperplane $H_{\lambda}:=\partial E\cap\Omega_{\lambda}$ is stable with respect to free-boundary variations in $\Omega_{\lambda}$, $\lambda\geq\lambda^{\ast}$, but our minimality result, Theorem 1.1, does not apply.
-We also remark that, despite $\Omega_{\lambda}$ not being a polyhedral cone, Theorem 1.1 appears to suggest that the dimensional bound for the singular set of the free-boundary proved in [8, Theorem 1.2 (1)] cannot be improved beyond the threshold n4n-4.
-In fact, when $n=4$ and $\lambda\leq 1/2\sqrt{2}$, the result above implies that $\partial E\cap\partial\Omega_{\lambda}=\{0\}$, and so $0\in\Sigma\cap\partial\Omega_{\lambda}$.
+$P(E;\Omega_{\lambda}\cap B_{R})\leq P(F;\Omega_{\lambda}\cap B_{R}).$ (1.3)
+In other words, if the opening angle $\alpha$ of the circular cone satisfies $\alpha\geq\bar{\alpha}(n):=\alpha_{\bar{\lambda}(n)}$, then $H_{\lambda}:=\partial E\cap\Omega_{\lambda}$ is a free-boundary area-minimizing surface inside the cone.
+We emphasize that the above theorem provides a definitive counterexample to the validity of a Vertex-skipping Theorem such as [28, Theorem 1.1] in higher dimensions, but only when $n\geq 4$.
+The case $n=3$ is a limiting situation where, in view of [27], provided $0<\lambda\leq\lambda^{\ast}(n)$, $H_{\lambda}$ is stable with respect to free-boundary variations in $\Omega_{\lambda}$, but our minimality result, Theorem 1.1, does not apply.
+We also remark that, despite $\Omega_{\lambda}$ not being a polyhedral cone, Theorem 1.1 appears to suggest that the dimensional bound for the singular set of the free-boundary proved in [10, Theorem 1.2 (1)] cannot be improved beyond the threshold n4n-4.
+In fact, when $n=4$ and $0<\lambda\leq 1/2\sqrt{2}$, the result above implies that $\partial E\cap\partial\Omega_{\lambda}=\{0\}$, and so $0\in\Sigma\cap\partial\Omega_{\lambda}$.
 The proof of Theorem 1.1 is based on a calibration argument.
 The calibration method was originally introduced by Bombieri, De Giorgi, and Giusti in their celebrated work [3] to prove the minimality of Simon’s cone in $\mathbb{R}^{8}$.
-Since then, several authors have contributed by proposing new versions of the original method [21, 10, 29, 5, 7] or by developing various generalizations to different contexts, such as the Steiner tree problem [27, 28, 33] and the Mumford-Shah functional [1, 30].
-Here, the calibration method is adapted to the free-boundary setting in the sense that we require the calibrating vector field $Z$ to be tangent to $\partial\Omega_{\lambda}$.
-This turns out to be crucial for carrying out the calibration argument, as it allows us to exclude the contribution coming from the area of $(E\Delta F)\cap\partial\Omega_{\lambda}$.
-We point out that the vector field $Z$ will not be defined in the entire closed cone $\overline{\Omega}_{\lambda}$, but only outside a $2$-dimensional subset of $\overline{\Omega}_{\lambda}$.
-This will force us to perform an approximation argument to apply the Divergence Theorem on $(E\Delta F)\cap\Omega_{\lambda}$, see Section 5.
-The existence of the vector field $Z$ will be shown in Section 4, and its construction will be carried out in two steps.
-Let $S_{\lambda}:=\partial\Omega_{\lambda}\setminus\{0\}$, and denote by $e_{1},\dots,e_{n},e_{t}$ the canonical basis of $\mathbb{R}^{n+1}$.
-We begin by calibrating the (n1)(n-1)-dimensional surface $S^{0}_{\lambda}:=H_{\lambda}\cap S_{\lambda}$ in $S_{\lambda}$; that is, we build a divergence-free vector field $Y$, defined on a suitable subset $S_{\lambda}^{\prime}$ of $S_{\lambda}$, such that $Y|_{S^{0}_{\lambda}}=e_{1}$ and $|Y|\leq 1$.
+Since then, several authors have contributed by proposing new versions of the original method [25, 12, 34, 30, 22, 4, 23, 6, 8] or by developing various generalizations to different contexts, such as the Steiner tree problem [32, 33, 39] and the Mumford-Shah functional [2, 1, 35].
+See [37, §6.5] for a more comprehensive overview on calibrations.
+Here, the calibration method is adapted to the free-boundary setting in the sense that we require the calibrating vector field to be tangent to $\partial\Omega_{\lambda}$.
+Our calibration argument, explained in detail in Section 5, proceeds roughly as follows.
+For $\lambda>0$, let $S_{\lambda}:=\partial\Omega_{\lambda}\setminus\{0\}$, and denote by $e_{1},\dots,e_{n},e_{t}$ the canonical basis of $\mathbb{R}^{n+1}$.
+First, we construct a divergence-free vector field $Z$, defined outside a $2$-dimensional subset of $\overline{\Omega}_{\lambda}$, such that $Z|_{H_{\lambda}}=e_{1}$, $|Z|\leq 1$, and $Z|_{S_{\lambda}}$ is tangent to $S_{\lambda}$.
+Second, we apply the Divergence Theorem to $Z$ on the region $(E\Delta F)\cap\Omega_{\lambda}$.
+On the one hand, the fact that $Z$ is tangent to $S_{\lambda}$ allows us to disregard the contribution arising from the area of $(E\Delta F)\cap\partial\Omega_{\lambda}$.
+On the other hand, the condition $\operatorname{div}Z=0$ enables us to compare $P(E;\Omega_{\lambda}\cap B_{R})$ with $P(F;\Omega_{\lambda}\cap B_{R})$.
+We stress that here, as remarked above, the vector field $Z$ is not defined on the entire closed cone $\overline{\Omega}_{\lambda}$, but only outside a $2$-dimensional subset of $\overline{\Omega}_{\lambda}$.
+This forces us to perform an approximation argument in order to apply the Divergence Theorem to $(E\Delta F)\cap\Omega_{\lambda}$ and infer the desired inequality (1.3); see Section 5.
+The existence of the vector field $Z$ will be exhibited in Section 4, and its construction will be carried out in two steps.
+We begin by calibrating the (n1)(n-1)-dimensional surface $S^{0}_{\lambda}:=H_{\lambda}\cap S_{\lambda}$ in $S_{\lambda}$; that is, we build a divergence-free vector field $Y$, defined on a suitable subset $S_{\lambda}^{\prime}$ of $S_{\lambda}$ and tangent to $S_{\lambda}$, such that $Y|_{S^{0}_{\lambda}}=e_{1}$ and $|Y|\leq 1$.
 In a slightly different formulation via differential forms, this was essentially done by Morgan in [36], using ideas from Lawlor’s PhD thesis [24].
-Second, we extend the calibration obtained in the previous step to (a suitable portion of) $\Omega_{\lambda}$ by vertical projection onto $S_{\lambda}$: we will show that, due to the structure of the metric on $S_{\lambda}$, the vector field $Z$ obtained in this way is, in particular, divergence-free in its domain.
+Their construction basically reduces the problem to finding a solution to a suitable differential inequality subject to appropriate initial conditions (cf.
+the proof of Theorem 4.2).
+Such a solution exists for $n\geq 4$ and $0<\lambda\leq\bar{\lambda}(n)$, thereby explaining the dimensional gap occurring for $n=3$.
+Second, we extend $Y$ to (a suitable portion of) $\Omega_{\lambda}$ by vertical projection onto $S_{\lambda}$: we will show that, due to the structure of the metric on $S_{\lambda}$, the vector field $Z$ obtained in this way is, in particular, divergence-free in its domain (cf.
+the proof of Theorem 4.1).
 1.1. Organization of the paper.
-In Section 2 we provide some preliminary notation and facts concerning differentiable manifolds (especially Riemannian ones) and sets of finite perimeter.
+In Section 2 we provide some preliminary notation and notions concerning differentiable manifolds (especially Riemannian ones) and sets of finite perimeter.
 In Section 3 we introduce further notation and perform some basic computations needed in the subsequent parts of the paper.
@@
 Given a vector field $X\in\mathfrak{X}(M)$, denote by $i_{X}\nu\in\mathcal{A}^{n-1}(M)$ the contraction between $X$ and the volume form $\nu$.
-We have the following technical Lemma (see [22, p.
+We have the following technical Lemma (see [26, p.
 52]).
@@
 $E^{(s)}:=\left\{(x,t)\in\mathbb{R}^{n+1}\,:\,\lim_{\rho\to 0^{+}}\dfrac{|E\cap B_{\rho}(x,t)|}{|B_{\rho}(x,t)|}=s\right\}.$
-If $E$ is a locally finite perimeter set, we indicate by $\partial^{\ast}E$ its reduced boundary and by $\nu_{E}$ the inner unit normal to its reduced boundary [26].
+If $E$ has locally finite perimeter, we indicate by $\partial^{\ast}E$ its reduced boundary and by $\nu_{E}$ the inner unit normal to $\partial^{\ast}E$ [31].
 Assume now that $E$ and $F$ have locally finite perimeter.
-We denote by $Q(E,F)^{\pm}$ the set
-$Q(E,F)^{\pm}:=\{(x,t)\in\partial^{\ast}E\cap\partial^{\ast}F\,:\,\nu_{E}(x,t)=\pm\nu_{F}(x,t)\}.$
-It is well known that also $E\cap F$ and $E\setminus F$ have locally finite perimeter and, up to ${\mathcal{H}}^{n}$-negligible sets, one has (see [26, Theorem 16.3])
+We define the set $Q(E,F)^{\pm}:=\{(x,t)\in\partial^{\ast}E\cap\partial^{\ast}F\,:\,\nu_{E}(x,t)=\pm\nu_{F}(x,t)\}.$ It is well known that $E\cap F$, $E\setminus F$ have also locally finite perimeter and, up to ${\mathcal{H}}^{n}$-negligible sets, one has (see [31, Theorem 16.3])
 $\partial^{\ast}(E\cap F)=\left(F^{(1)}\cap\partial^{\ast}E\right)\cup\left(E^{(1)}\cap\partial^{\ast}F\right)\cup Q(E,F)^{+}$ (2.6)
@@
 $\partial_{i}=e_{i}+A_{i}\,e_{t}\qquad\text{where}\qquad A_{i}:=\lambda\,\dfrac{x_{i}}{|x|}\,.$
-Standard properties of the determinant guarantee that $\det g=|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}$. Define the multi-index $\beta=(1,\dots,n-1)$. We have
-$\partial_{1}\wedge\dots\wedge\partial_{n}$ $=e_{\beta}\wedge e_{n}+A_{n}\,e_{\beta}\wedge e_{t}+\sum_{i=1}^{n-1}(-1)^{n-1-i}A_{i}\,e^{i}_{\beta}\wedge e_{t}\wedge e_{n}.$ (3.7)
-Now the wedge products appearing in (3.7) are orthonormal each other, hence
-$|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}=1+\sum_{i=1}^{n}A_{i}^{2}=1+\lambda^{2},$
-and this proves (3.4).
+Standard properties of the determinant guarantee that $\det g=|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}$. Define the multi-index $\beta=(1,\dots,n)$. We have
+$\partial_{1}\wedge\dots\wedge\partial_{n}$ $=e_{\beta}+\sum_{i=1}^{n}(-1)^{n-i}A_{i}\,e^{i}_{\beta}\wedge e_{t}.$ (3.7)
+Now we observe that the wedge products appearing in (3.7) are orthonormal each other, hence $|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}=1+\sum_{i=1}^{n}A_{i}^{2}=1+\lambda^{2},$ and this proves (3.4).
 Let us now show the validity of (3.5). We start observing that, by the definitions of $\rho$ and $\theta$,
@@
 Denote by $\tilde{g}_{1}$ the $(n-1)\times(n-1)$ submatrix of $\tilde{g}$ obtained by cancelling the first row and the first column of $\tilde{g}$.
-By the definition of the metric on $\bigwedge^{k}T^{\ast}M$ given in (2.2) (in this case k=n2k=n-2) and the properties of the adjugate matrix (see [34, p.
+By the definition of the metric on $\bigwedge^{k}T^{\ast}M$ given in (2.2) (in this case k=n2k=n-2) and the properties of the adjugate matrix (see [40, p.
 73]), we have
@@
 $h(0)=1,\qquad h^{\prime}(0)=0\qquad\text{and}\qquad|\omega_{h}(x)|_{g}\leq 1\quad\text{for all $x\in M^{\prime}$.}$ (4.5)
-Applying the Cauchy-Schwartz inequality for wedge products (see [9, pag. 32], and note that $1$-forms are always simple), by (4), we obtain
+Applying the Cauchy-Schwartz inequality for wedge products (see [11, pag. 32], and note that $1$-forms are always simple), by (4), we obtain
 $|\omega_{h}(x)|_{g}\leq\left|\left(h(u)-\dfrac{x_{1}}{(n-1)r}h^{\prime}(u)\right)dr+\dfrac{h^{\prime}(u)}{n-1}dx_{1}\right|_{g}\cdot\,\,\left|\dfrac{n-1}{r}\psi_{0}(x)\right|_{g}\,.$
@@
 $P(E;(F^{(0)}\cup F^{(1)})\cap\Omega_{\lambda}^{\varepsilon}\cap B_{1})\leq P(F;(\bar{E}^{c}\cup E)\cap\Omega_{\lambda}^{\varepsilon}\cap B_{1})+C(n)\,\varepsilon^{n-2}.$
-By applying Federer’s Theorem [26, Theorem 16.2], we then get
+By applying Federer’s Theorem [31, Theorem 16.2], we then get
 $P(E;\Omega_{\lambda}^{\varepsilon}\cap B_{1})\leq P(F;\Omega_{\lambda}^{\varepsilon}\cap B_{1})+C(n)\,\varepsilon^{n-2}.$